Solve: 2^x + x = 5 , x is real
Answer by ikleyn(53763)

.

This equation can not be solved algebraically.

It can be solved approximately, using numerical methods.

Go to web-site www.desmos.com/calculator,
print there y = 2^x + x - 5 and get the plot.

Click where the plot intersects x-axis. You will see the number 1.716.

It is an approximate solution and your answer.

Solved, with explanations.


/////////////////


        Also,  there are supporting thoughts that may help you to feel more confidence.


Notice that left side of the equation is the function,  which monotonically increases.

At  x= 1,   left side is   2 + 1 = 3.

At  x= 2,   left side is   4 + 2 = 6.

It tells you that the solution does exist,  that there is only one solution
(= the solution is unique),  and that the solution is between  1  and  2.